At what temperature will 0.7 mole of chlorine exert a pressure of 1.18 atm at a volume of 0.75L?


1
Expert's answer
2021-02-23T04:36:13-0500

find the formula:

"PV = nRT"

"R =\frac{0.0821 atm.L}{mol.K}"

Volume = 0.75 L

Now we will put the values in formula.

"T = \frac{PV}{nR}"

"T = \frac{1.18 atm × 0.75 L}{ 0.07 mol ×0.0821 atm.L/mol.K}"

T = 154 K